# Einstein Telescope: Detection of Binary Black Hole Gravitational Wave Signals Using Deep Learning

### v3

@@ Abstract
Expanding upon our prior work (Alhassan et al., 2022), where we evaluated the performance of single sub-detector data (SSDD) from the Einstein Telescope (ET) for binary black hole (BBH) detection using deep learning (DL). In this study, we explore the detection efficiency of BBHs using data combined from all three proposed sub-detectors of ET (TSDCD), employing five different lower frequency cutoffs ($F_{\text{low}}$): 5 Hz, 10 Hz, 15 Hz, 20 Hz, and 30 Hz, while maintaining the same match-filter Signal-to-Noise Ratio ($MSNR$) ranges as in our previous work: 4-5, 5-6, 6-7, 7-8, and >8. The Deep Residual Neural Network model (ResNet) was trained and evaluated for the detection of BBH gravitational wave signals using both TSDCD and SSDD. Compared to SSDD, the detection accuracy from TSDCD has shown substantial improvements, increasing from $60\%$, $60.5\%$, $84.5\%$, $94.5\%$ to $78.5\%$, $84\%$, $99.5\%$, $100\%$, and $100\%$ for sources with $MSNR$ of 4-5, 5-6, 6-7, 7-8, and >8, respectively. In a qualitative evaluation, the ResNet model detected sources at 86.601 Gpc, with an averaged $MSNR$ of 3.9 (averaged across the three sub-detectors) and a chirp mass of 13.632 at 5 Hz. The results demonstrate a notable accuracy improvement for lower $MSNR$ ranges (4-5, 5-6, 6-7) by $18.5\%$, $24.5\%$, and $13\%$, respectively, and by $5.5\%$ and $1.5\%$ for higher $MSNR$ ranges (7-8 and >8). TSDCD proves suitable for near-real-time detection and can benefit from a more powerful setup.
@@ 1 Introduction
This is a continuation of our previous work (Alhassan et al., 2022) (WTM1 hereafter referred to as) on the detection of binary black holes (BBHs) gravitational wave (GWs) signals from the Einstein Telescope (ET) using deep learning (DL). ET (Punturo et al., 2010; Abernathy et al., 2011; Maggiore et al., 2020) is designed [^1] See https://www.et-gw.eu/index.php/relevant-et-documents for relevant documents on the ET design. to be an underground GWs detector, where the seismic noise is much lower, and hence the level of Newtonian noise. ET as shown in Figure 1, will consist of three nested arms (hereafter referred to as sub-detectors) with 10 Km long each, in an equilateral triangle shape, with arm-opening angles of $60^{\circ}$ (Hild et al., 2009; Branchesi et al., 2023). Compared to 3 km length for advanced Virgo (Acernese et al., 2014) and KAGRA (Aso et al., 2013; Akutsu et al., 2020), and 4 km for advanced LIGO (Collaboration et al., 2015), the ET 10 km V-shaped sub-detectors length will significantly improve its sensitivity allowing the observation of GWs at lower frequency that reach $\approx 2$ Hz (Hild et al., 2011). Each sub-detector of ET will consist of two interferometers, one optimised for low-frequency and one for high-frequency sensitivity. In this work, it is assumed that each sub-detector has a single interferometer.
@@
The application examples of deep learning in astrophysics, and more specifically in GWs, were shown in WTM1. Notably, more work has been presented since then, highlighting the significant interest in this promising technology for addressing the considerable challenge posed by the vast amount of data expected from current and upcoming detectors. The Spy project (Glanzer et al., 2023) applied machine learning for glitch classification using data up to the end of the third observing run of LIGO. In this analysis, 233,981 glitches from LIGO Hanford and 379,805 glitches from LIGO Livingston were classified into morphological classes. A related study employed generative adversarial networks (GANs) to simulate hundreds of synthetic images representing the 22 most common types of glitches observed in the LIGO, KAGRA, and Virgo detectors. A neural network model was then used to classify the simulated glitches, achieving an average accuracy of $99.0\%$ (Powell et al., 2023). Accelerating Artificial Intelligence models for rapid GW detection was recently conducted by Chaturvedi et al. (2022). In thier work, the ThetaGPU supercomputer was leveraged with the NVIDIA TensorRT-optimized AI ensemble to process a month of advanced LIGO data; the process took 50 seconds. In Yan et al. (2022), a novel model called a gravitational wave squeeze-and-excitation shrinkage network (GW-SESNet) was inspired by attention mechanisms, a mechanism in deep learning used to highlight important features, for the classification of real signals, noise, and glitches. GW-SESNet utilized the coherent signal-to-noise ratio (SNR) from multiple detectors and achieved recall values of 0.84, 0.98, and 0.98 for noise, real signals, and glitches, respectively. The role of advanced machine learning tools in the coming decades, either independently or in tandem with traditional data analysis techniques, may become crucial considering the vast amount of the data expected (Shah et al., 2023).
@@
The structure of this work goes as follow: in Section 2 we describe the method and tools were used for the generation of TSDCD and the prepossessing pipeline employed for spectrogams preparation. Section 3 discusses Deep Residual Neural Networks, and the training and evaluation processes. Evaluating ResNet model on the TSDCD is presented in section 4 and 5. Finally, a summary of this work and a future work plans is presented in Section 6.
@@ 2 ET data simulation
The strain in each sub-detector of ET $h(t)_{E\alpha}$ is given as :
$h(t)_{E\alpha}=\frac{1}{4}F_{\alpha}^{+}(\theta,\phi,\psi)h_{+}(t)+F_{\alpha}^{\times}(\theta,\phi,\psi)h_{\times}(t),$ (1)
@@
$a(t)$ $=\frac{1}{16}\sin 2\gamma(3-\cos 2\lambda)(3-\cos 2\delta)\cos[2(\alpha-\phi_{r}-\Omega_{r}t)]$
$\quad-\frac{1}{4}\cos 2\gamma\sin\lambda(3-\cos 2\delta)\sin[2(\alpha-\phi_{r}-\Omega_{r}t)]$
$\quad+\frac{1}{4}\sin 2\gamma\sin 2\lambda\sin 2\delta\cos[\alpha-\phi_{r}-\Omega_{r}t]$
$\quad-\frac{1}{2}\cos 2\gamma\cos\lambda\sin 2\delta\sin[\alpha-\phi_{r}-\Omega_{r}t]$
$\quad+\frac{3}{4}\sin 2\gamma\cos 2\lambda\cos 2\delta,$ (4)
$b(t)$ $=\cos 2\gamma\sin\lambda\sin\delta\cos[2(\alpha-\phi_{r}-\Omega_{r}t)]$
$\quad+\frac{1}{4}\sin 2\gamma(3-\cos 2\lambda)\sin\delta\sin[2(\alpha-\phi_{r}-\Omega_{r}t)]$
$\quad+\cos 2\gamma\cos\lambda\cos\delta\cos[\alpha-\phi_{r}-\Omega_{r}t]$
$\quad+\frac{1}{2}\sin 2\gamma\sin 2\lambda\cos\delta\sin[\alpha-\phi_{r}-\Omega_{r}t],$ (5)
@@
$h_{+}(t)$ $=-\frac{1+\cos^{2}\iota}{2}\left(\frac{GMc^{2}D_{L}}{2(tc-t)}\right)^{1/4}\times\cos\left[2\Phi_{c}+2\Phi(t-tc;M,\mu)\right],$ (6)
$h_{\times}(t)$ $=-\cos\iota\left(\frac{GMc^{2}D_{L}}{2(tc-t)}\right)^{1/4}\times\sin\left[2\Phi_{c}+2\Phi(t-tc;M,\mu)\right],$ (7)
@@ 2.1 Simulation
Assuming the standard $\Lambda$CDM cosmology (Planck Collaboration et al., 2016), where the Hubble constant $H_{0}=67.74\,\text{km s}^{-1}\,\text{Mpc}^{-1}$, energy density parameter $\Omega_{\Lambda}=0.6910$, and matter density parameter $\Omega_{m}=0.3075$. We convert the redshift $z$ to $D_{L}$ using the analytical approximation provided by Adachi & Kasai (2012) for the $D_{L}$ between the BBH source and the detector, given as:
$D_{L}=\left(1+z\right)cH_{0}\int_{0}^{z}\frac{dz}{\sqrt{\Omega_{\Lambda}+\Omega_{m}(1+z)^{3}}}$ (9)
where $c$ is the speed of light in the vacuum.
We utilized the latest time-domain phenomenological higher-order mode model IMRPhenomD (Pürrer et al., 2023), which is part of LALSuite (LIGO Scientific Collaboration et al., 2018), employing its Python interface (Wette, 2020). IMRPhenomD is designed for generating GW signals from merging BBH with non-precessing spins. For each sub-detector, using the parameters above, we individually generated a BBH GW signal ($h(t)_{E1},h(t)_{E2},h(t)_{E3}$), then added them into the simulated Gaussian detector noise $n$, assuming an additive, Gaussian and stationary noise model as illustrated in Figure 2. This process results in two datasets: The data $d$ (if the signal $h$ is present) and $d_{n}$ (if the data contains noise only), can be written as follow:
$d(t)$ $=h(t)_{E1}+h(t)_{E2}+h(t)_{E3}+n(t)$ (10)
$d_{n}(t)$ $=n(t)$ (11)
Figure 2: A 5-second segment of simulated data for all ET sub-detectors. The gray line represents detector noise $n(t)$. The blue, green, and red lines represent the injected BBH signals from E1, E2, and E3, respectively. The injected BBH source has an MSNR in the range $7-8$ with component masses $m_{1}=20\,M_{\odot}$ and $m_{2}=25\,M_{\odot}$, and $F_{\text{low}}$ of 30 Hz.
The improvement in SNR through the utilization of data from all sub-detectors of ET combined can be assessed by calculating the coherent SNR. Coherent SNR $\text{SNR}_{\text{coh}}$ enhances the observed signal’s SNR by incorporating the coherent combination of multiple detectors (Macleod et al., 2016). To quantify the improvement in MSNR resulting from the combination of each sub-detector, we calculated $\text{SNR}_{\text{coh}}$ using the following formula:
@@ 2.2 Spectrograms
The Short Time Fourier Transform (STFT) (Yin et al., 2013) was used to generate our TSDCD spectrogram samples ($S_{E123}$), as shown in Figure 4. Three spectrograms ($S_{E1}$, $S_{E2}$, and $S_{E3}$) were produced for each source in each sub-detector ($E1$, $E2$, and $E3$), and then stacked to form an RGB image $S_{E123}$, represented as:
@@
Similarly, to generate only-noise spectrograms, the following steps were taken: 1) A Gaussian noise with a color matching the power spectrum density of ET, sampled at ET’s frequency, was generated. 2) Three random segments of 5 seconds each were selected. 3) STFT was performed on each segment to produce a spectrogram. 4) All three only-noise spectrograms were stacked together.
@@ 3 ResNet: Deep Residual Neural Networks
One of the great advantages of ResNet network, is its ability to overcome the vanishing gradient problem, which considered as one of the challenging issues in training deep and complex neural networks models (Hu et al., 2018; Roodschild et al., 2020). This was achieved by introducing residual blocks, which enable the training of hundreds of layers (Goceri, 2019a, b; Hasan et al., 2021). These blocks use an additive attribute to merge previous and future layers. ResNet is composed of several residual blocks. Each residual block is formed by adding a residual connection every two layers of conventional convolution. The key concept behind the residual connection, which is considered a pivotal breakthrough in deep learning, is to learn a residual function capable of representing the difference between the input $X$ and the target $\mathcal{F}(X)$ (the desired mapping learned by the layer). A residual mapping $\mathcal{R}(X)$ is learned as follows:
@@
Then the sum of the residual mapping and the input is the output of the layer ($Y$) and can be calculated as:
@@
Due to this connection, the original output will be superimposed before sending it to the future layer. And hence, mitigate and prevent any gradient explosion or disappearance when using the backpropagation (Munro, 2010, for details about backpropagation please see), and allow for training of very deep networks. As previously mentioned in WTM1, ResNet is available in various architectures, distinguished by the number of layers. This includes ResNet-50 and ResNet-101, which comprise 50 and 101 convolutional layers, respectively. We adopt the ResNet-101 architecture (hereafter referred to as ResNet) in this work, as was previously done in WTM1.

### v6

# PyMerger: Detecting Binary Black Hole mergers from Einstein Telescope Using Deep Learning

@@ Abstract
We present PyMerger, a Python tool for detecting binary black hole (BBH) mergers from the Einstein Telescope (ET), based on a Deep Residual Neural Network model (ResNet). ResNet was trained on data combined from all three proposed sub-detectors of ET (TSDCD) to detect BBH mergers. Five different lower frequency cutoffs ($F_{\text{low}}$): 5 Hz, 10 Hz, 15 Hz, 20 Hz, and 30 Hz, with match-filter Signal-to-Noise Ratio ($MSNR$) ranges: 4-5, 5-6, 6-7, 7-8, and >8, were employed in the data simulation. Compared to previous work that utilized data from single sub-detector data (SSDD), the detection accuracy from TSDCD has shown substantial improvements, increasing from $60\%$, $60.5\%$, $84.5\%$, $94.5\%$ to $78.5\%$, $84\%$, $99.5\%$, $100\%$, and $100\%$ for sources with $MSNR$ of 4-5, 5-6, 6-7, 7-8, and >8, respectively. The ResNet model was evaluated on the first Einstein Telescope mock Data Challenge (ET-MDC1) dataset, where the model demonstrated strong performance in detecting BBH mergers, identifying 5,566 out of 6,578 BBH events, with optimal SNR starting from 1.2, and a minimum and maximum $D_{L}$ of 0.5 Gpc and 148.95 Gpc, respectively. Despite being trained only on BBH mergers without overlapping sources, the model achieved high BBH detection rates. Notably, even though the model was not trained on BNS and BHNS mergers, it successfully detected 11,477 BNS and 323 BHNS mergers in ET-MDC1, with optimal SNR starting from 0.2 and 1, respectively, indicating its potential for broader applicability.
@@ 1 Introduction
This is a continuation of our previous work (Alhassan et al., 2022) (hereafter referred to as WTM1) on the detection of binary black holes (BBHs) gravitational wave (GWs) signals from the Einstein Telescope (ET) using deep learning (DL). ET (Punturo et al., 2010; Abernathy et al., 2011; Maggiore et al., 2020) is designed [^1] See https://www.et-gw.eu/index.php/relevant-et-documents for relevant documents on the ET design. to be an underground GWs detector, where the seismic noise is much lower, and hence the level of Newtonian noise. ET as shown in Figure 1, will consist of three nested arms (hereafter referred to as sub-detectors) with 10 Km long each, in an equilateral triangle shape, with arm-opening angles of $60^{\circ}$ (Hild et al., 2009; Branchesi et al., 2023). Compared to 3 km length for advanced Virgo (Acernese et al., 2014) and KAGRA (Aso et al., 2013; Akutsu et al., 2020), and 4 km for advanced LIGO (Collaboration et al., 2015), the ET 10 km V-shaped sub-detectors length will significantly improve its sensitivity allowing the observation of GWs at lower frequency that reach $\approx 2$ Hz (Hild et al., 2011). Each sub-detector of ET will consist of two interferometers, one optimised for low-frequency and one for high-frequency sensitivity. In this work, it is assumed that each sub-detector has a single interferometer.
@@
The application examples of deep learning in astrophysics, and more specifically in GWs, were shown in WTM1. Notably, more work has been presented since then, highlighting the significant interest in this promising technology for addressing the considerable challenge posed by the vast amount of data expected from current and upcoming detectors. The Spy project (Glanzer et al., 2023) applied machine learning for glitch classification using data up to the end of the third observing run of LIGO. In this analysis, 233,981 glitches from LIGO Hanford and 379,805 glitches from LIGO Livingston were classified into morphological classes. A related study employed generative adversarial networks (GANs) to simulate hundreds of synthetic images representing the 22 most common types of glitches observed in the LIGO, KAGRA, and Virgo detectors. A neural network model was then used to classify the simulated glitches, achieving an average accuracy of $99.0\%$ (Powell et al., 2023). Accelerating Artificial Intelligence models for rapid GW detection was recently conducted by Chaturvedi et al. (2022). In their work, the ThetaGPU supercomputer was leveraged with the NVIDIA TensorRT-optimized AI ensemble to process a month of advanced LIGO data; the process took 50 seconds. In Yan et al. (2022), a novel model called a gravitational wave squeeze-and-excitation shrinkage network (GW-SESNet) was inspired by attention mechanisms, a mechanism in deep learning used to highlight important features, for the classification of real signals, noise, and glitches. GW-SESNet utilized the coherent signal-to-noise ratio (SNR) from multiple detectors and achieved recall values of 0.84, 0.98, and 0.98 for noise, real signals, and glitches, respectively. The role of advanced machine learning tools in the coming decades, either independently or in tandem with traditional data analysis techniques, may become crucial considering the vast amount of the data expected (Shah et al., 2023).
@@
The structure of this work goes as follow: in Section 2 we describe the method and tools were used for the generation of TSDCD and the prepossessing pipeline employed for spectrogams preparation. Section 3 discusses Deep Residual Neural Networks, and the training and evaluation processes. Evaluating ResNet model on noncontinuous and continuous TSDCD is presented in section 4 and 5. Evaluation on the first ET mock data challenge is described in section 6. Description of PyMerger software is presented in section 7. Finally, a summary of this work and a future work plans is presented in Section 8.
@@ 2 ET data simulation
The strain in each sub-detector of ET $h_{E\alpha}(t)$ is given as :
$h_{E\alpha}(t)=F_{\alpha}^{+}(\theta,\phi,\psi)h_{+}(t)+F_{\alpha}^{\times}(\theta,\phi,\psi)h_{\times}(t),$ (1)
@@
$\begin{split}a(t)={}&\frac{1}{16}\sin 2\gamma(3-\cos 2\lambda)(3-\cos 2\delta)\cos[2(\alpha-\phi_{r}-\Omega_{r}t)]\\ &-\frac{1}{4}\cos 2\gamma\sin\lambda(3-\cos 2\delta)\sin[2(\alpha-\phi_{r}-\Omega_{r}t)]\\ &+\frac{1}{4}\sin 2\gamma\sin 2\lambda\sin 2\delta\cos[\alpha-\phi_{r}-\Omega_{r}t]\\ &-\frac{1}{2}\cos 2\gamma\cos\lambda\sin 2\delta\sin[\alpha-\phi_{r}-\Omega_{r}t]\\ &+\frac{3}{4}\sin 2\gamma\cos 2\lambda\cos 2\delta,\end{split}$ (4)
$\begin{split}b(t)={}&\cos 2\gamma\sin\lambda\sin\delta\cos[2(\alpha-\phi_{r}-\Omega_{r}t)]\\ &+\frac{1}{4}\sin 2\gamma(3-\cos 2\lambda)\sin\delta\sin[2(\alpha-\phi_{r}-\Omega_{r}t)]\\ &+\cos 2\gamma\cos\lambda\cos\delta\cos[\alpha-\phi_{r}-\Omega_{r}t]\\ &+\frac{1}{2}\sin 2\gamma\sin 2\lambda\cos\delta\sin[\alpha-\phi_{r}-\Omega_{r}t],\end{split}$ (5)
@@
$h_{+}(t)=-\frac{1+\cos^{2}\iota}{2}\left(\frac{G\mathcal{M}}{c^{2}D_{L}}\right)\left(\frac{t_{c}-t}{G\mathcal{M}/c^{3}}\right)^{-1/4}\cos\left[2\Phi_{c}+2\Phi(t-t_{c};M,\mu)\right]$ (6)
$h_{\times}(t)=-\cos\iota\left(\frac{G\mathcal{M}}{c^{2}D_{L}}\right)\left(\frac{t_{c}-t}{G\mathcal{M}/c^{3}}\right)^{-1/4}\sin\left[2\Phi_{c}+2\Phi(t-t_{c};M,\mu)\right]$ (7)
@@ 2.1 Simulation
Assuming the standard $\Lambda$CDM cosmology (Planck Collaboration et al., 2016), where the Hubble constant $H_{0}=67.74\,\text{km s}^{-1}\,\text{Mpc}^{-1}$, energy density parameter $\Omega_{\Lambda}=0.6910$, and matter density parameter $\Omega_{m}=0.3075$. We convert the redshift $z$ to $D_{L}$ using the analytical approximation provided by Adachi & Kasai (2012) for the $D_{L}$ between the BBH source and the detector, given by:
$d_{L}(z,\Omega_{m})=\frac{2c}{H_{0}\sqrt{\Omega_{m}}}(1+z)\left\{\Phi\left(x(0,\Omega_{m})\right)-\frac{1}{\sqrt{1+z}}\Phi\left(x(z,\Omega_{m})\right)\right\},$ (9)
where $c$ is the speed of light in vacuum, and $\Phi(x)$ is a rational fitting function given by:
$\Phi(x)=\frac{1+1.320x+0.4415x^{2}+0.02656x^{3}}{1+1.392x+0.5121x^{2}+0.03944x^{3}},$ (10)
and $x(z,\Omega_{m})$ is defined as:
$x(z,\Omega_{m})=\frac{1-\Omega_{m}}{\Omega_{m}}\cdot\frac{1}{(1+z)^{3}}.$ (11)
We utilized the frequency-domain phenomenological non spinning BBH model IMRPhenomD (Husa et al., 2016), which is part of LALSuite (LIGO Scientific Collaboration et al., 2018), employing its Python interface (Wette, 2020). IMRPhenomD is designed for generating GW signals from merging BBH with non-precessing spins. For each sub-detector, using the parameters above, we individually generated a BBH GW signal ($h(t)_{E1},h(t)_{E2},h(t)_{E3}$), then added them into the simulated Gaussian detector noise $n$, assuming an additive, Gaussian and stationary noise model. This process results in tow datasets: The data $d_{1,2,3}$ (if the signal $h$ is present) and $d_{n}$ (if the data contains noise only), can be written as follow:
$d_{1}(t)$ $=h_{E1}(t)+n_{1}(t)$ (12)
$d_{2}(t)$ $=h_{E2}(t)+n_{2}(t)$ (13)
$d_{3}(t)$ $=h_{E3}(t)+n_{3}(t)$ (14)
$d_{n}(t)$ $=n_{1}(t)+n_{2}(t)+n_{3}(t)$ (15)
Figure 2: A 5-second segment of simulated noise from a single sub-detector is overlapped with BBH GW signals. The gray line represents a noise from a single sub-detector noise $n_{1}(t)$. The blue, green, and red lines represent the BBH signals from E1, E2, and E3, respectively. The simulated BBH signal has an MSNR in the range $7-8$ with component masses $m_{1}=20\,M_{\odot}$ and $m_{2}=25\,M_{\odot}$, and $F_{\text{low}}$ of 30 Hz.
To illustrate how the GW signal would look and compare it to the detector’s noise, Figure 2 shows the difference in phase between signals from the three sub-detectors and the noise from a single sub-detector in the background. The improvement in SNR through the utilization of data from all sub-detectors of ET combined can be assessed by calculating the coherent SNR. Coherent SNR $\text{SNR}_{\text{coh}}$ enhances the observed signal’s SNR by incorporating the coherent combination of multiple detectors (Macleod et al., 2016). To quantify the improvement in MSNR resulting from the combination of each sub-detector, we calculated $\text{SNR}_{\text{coh}}$ using the following formula:
@@ 2.2 Spectrograms
The Short Time Fourier Transform (STFT) (Yin et al., 2013) was used to generate our TSDCD spectrogram samples ($S_{E123}$), as shown in Figure 4. Three spectrograms ($S_{E1}$, $S_{E2}$, and $S_{E3}$) were produced for each injected segemnets, $d_{1,2,3}(t)$ from each sub-detector ($E1$, $E2$, and $E3$) respectively, and then stacked to form an RGB image $S_{E123}$, represented as:
@@
Similarly, to generate only-noise spectrograms, the following steps were taken: 1) A Gaussian noise with a color matching the power spectrum density of ET, sampled at ET’s frequency, was generated for each sub-detector. 2) Three random segments of 5 seconds each were selected from each sub-detector’s noise, $n_{1}(t)$, $n_{2}(t)$, and $n_{3}(t)$. 3) STFT was performed on each segment to produce a spectrogram. 4) All three only-noise spectrograms were stacked together.
@@ 3 ResNet: Deep Residual Neural Networks
One of the great advantages of ResNet network, is its ability to overcome the vanishing gradient problem, which considered as one of the challenging issues in training deep and complex neural networks models (Hu et al., 2018; Roodschild et al., 2020). This was achieved by introducing residual blocks, which enable the training of hundreds of layers (Goceri, 2019a, b; Hasan et al., 2021). These blocks use an additive attribute to merge previous and future layers. ResNet is composed of several residual blocks. Each residual block is formed by adding a residual connection every two layers of conventional convolution. The key concept behind the residual connection, considered a pivotal breakthrough in deep learning, is to learn a residual function capable of representing the difference between the input $X$ and the target $\mathcal{F}(X)$ (the desired mapping learned by the layer). A residual mapping $\mathcal{R}(X)$ is learned as follows:
@@
Then, the sum of the residual mapping and the input is the output of the layer ($Y$), and can be calculated as:
@@
In the ResNet architecture, $Y$ is equivalent to $2\mathcal{F}(X)-X$, reflecting the residual connection $Y=f(x)+f(x)-x$, where $f(x)$ represents the output of a series of trainable layers.
Due to this connection, the original output will be superimposed before sending it to the future layer. And hence, mitigate and prevent any gradient explosion or disappearance when using the backpropagation (Munro, 2010, for details about backpropagation please see), and allow for training of very deep networks. More details about ResNet can be found in appendix A. As previously mentioned in WTM1, ResNet is available in various architectures, distinguished by the number of layers. This includes ResNet-50 and ResNet-101, which comprise 50 and 101 convolutional layers, respectively. We adopt the ResNet-101 architecture (hereafter referred to as ResNet) in this work, as was previously done in WTM1.